# Optimal Rank-Metric Codes with Rank-Locality from Drinfeld Modules

### v1

@@ 5.3. Proof of the main theorem
Let $m\geq 4$ be an integer such that
@@
Then there exists a monic irreducible polynomial $P\in{\mathbb{F}}_{q}[T]$ of degree $m$ such that $P\equiv 1\pmod{h}$.
@@
For $a=1$ and $h=\prod^{\ell}_{i=1}(T^{R}-a_{i})$ in Theorem 5.1 we know that
|#$\left|\#S_{m}(1,h)-\frac{q^{m}}{m\Phi(h)}\right|\leq 3\frac{q^{\frac{m}{2}}}{m}T(m)+2\deg(h)\frac{q^{\frac{m}{2}}}{m}.$
@@
#$\#S_{m}(1,h)\geq\frac{q^{m}}{m\Phi(h)}-6\frac{q^{\frac{m}{2}}}{\sqrt{m}}-2q^{\frac{m}{2}}.$
Bounding now $\Phi(h)$ simply with $q^{\ell R}$ and observing that we want to estimate $m$ for which #$\#S_{m}(1,h)$ is at least $1$, it is sufficient to check when
@@
Finally, we are ready to prove Theorem 1.1.
Proof of Theorem 1.1.
With the framework we constructed, the proof of our result becomes immediate. First, since we have
@@
Lemma 5.2 guarantees the existence of an irreducible polynomial $P$ of degree $m$ that is congruent to $1$ modulo $h$. This allows for the construction of Section 3 to happen, as long as the conditions C1 to C4 are satisfied. Notice that C1,C2,C3 are directly satisfied by the hypothesis of Theorem 1.1, and C4 is forced by the stronger condition arising from Lemma 5.2 forcing the existence of $P$. This guarantees that the construction applies. Now derive the wanted parameters from Theorem 3.4. ∎
6. Acknowledgements
We would like to thank Federico Pellarin for interesting conversations on the topics of this paper. This work was supported by the National Science Foundation under Grant No 2338424.
Table 1.
[…]
5 3 3 $T^{9}+4T^{6}+T^{3}+4$ 10 $T+2$ $T^{10}+2T^{9}+4T^{7}+3T^{6}+T^{4}+2T^{3}+4T+4$
11 $T^{2}+T$ $T^{11}+T^{10}+4T^{8}+4T^{7}+T^{5}+T^{4}+4T^{2}+4T+1$
12 $T^{3}+2$ $T^{12}+T^{9}+4T^{6}+T^{3}+4$
13 $T^{4}+2T+4$ $T^{13}+T^{10}+4T^{9}+4T^{7}+T^{6}+T^{4}+4T^{3}+3T+2$
14 $T^{5}+4T$ $T^{14}+4T^{11}+4T^{10}+T^{8}+T^{7}+4T^{5}+4T^{4}+T+1$
$(a_{1},a_{2},a_{3})=(1,2,3)$ 15 $T^{6}+T+2$ $T^{15}+4T^{12}+T^{10}+3T^{9}+4T^{7}+2T^{6}+T^{4}+2T^{3}+4T+4$
16 $T^{7}+3$ $T^{16}+4T^{13}+T^{10}+3T^{9}+4T^{7}+2T^{6}+3T^{3}+3$
17 $T^{8}$ $T^{17}+4T^{14}+T^{11}+4T^{8}+1$
7 3 2 $T^{6}+4T^{3}+2$ 7 $T$ $T^{7}+4T^{4}+2T+1$
8 $T^{2}+1$ $T^{8}+T^{6}+4T^{5}+4T^{3}+2T^{2}+3$
9 $T^{3}+1$ $T^{9}+5T^{6}+6T^{3}+3$
$(a_{1},a_{2})=(1,2)$ 10 $T^{4}+T+5$ $T^{10}+5T^{7}+5T^{6}+6T^{4}+6T^{3}+2T+4$
11 $T^{5}+6$ $T^{11}+4T^{8}+6T^{6}+2T^{5}+3T^{3}+6$
8 $T^{2}$ $T^{8}+(\alpha+1)T^{5}+\alpha T^{2}+1$
$(a_{1},a_{2})=(1,\alpha)$ 10 $T^{4}$ $T^{10}+(\alpha+1)T^{7}+\alpha T^{4}+1$
11 $T^{5}$ $T^{11}+(\alpha+1)T^{8}+\alpha T^{5}+1$

### v2

@@ 6.2. Existence of an infinite family
Set $H=\prod^{\ell}_{i=1}(T^{R}-a_{i})$. Let $m\geq 4$ be an integer such that
@@
Then there exists a monic irreducible polynomial $\mathfrak{q}\in{\mathbb{F}}_{q}[T]$ of degree $m$ such that $\mathfrak{q}\equiv 1\pmod{h}$.
@@
For $a=1$ and in Theorem 6.8 we know that
$\left|\#S_{m}(1,h)-\frac{q^{m}}{m\Phi(h)}\right|\leq 3\frac{q^{\frac{m}{2}}}{m}T(m)+2\deg(h)\frac{q^{\frac{m}{2}}}{m}.$
@@
$S_{m}(1,h)\geq\frac{q^{m}}{m\Phi(h)}-6\frac{q^{\frac{m}{2}}}{\sqrt{m}}-2q^{\frac{m}{2}}.$
Bounding now $\Phi(h)$ simply with $q^{\ell R}$ and observing that we want to estimate $m$ for which $S_{m}(1,h)$ is at least $1$, it is sufficient to check when
@@
Finally, we are ready to prove the existence of an infinite family of our codes.
Theorem 6.10.
@@
$2\leq\delta<R$, $s+1\leq\ell<q$. Then, there exists a rank-metric code in ${\mathbb{F}}_{q}^{m\times\ell R}$, rank-locality $(r,\delta)$, $R=r+\delta-1$, and ${\mathbb{F}}_{q}$-dimension equal to $m(s+1)r$. Moreover, this code has rank distance $\ell R-Rs-r+1$, thus it is optimal with respect to the bound in Proposition 2.22.
Proof.
Employ the construction of subsection 4.2 observing that with the hypothesis of the Theorem the parameters verify conditions C0, C1, C2 (and therefore (1) of subsection 4.2). Now, since we have
Lemma 6.9 guarantees the existence of an irreducible polynomial $\mathfrak{q}$ of degree $m$ that is congruent to $1$ modulo $H$, which verifies (3) in the construction of subsection 4.2. This allows for the construction of subsection 4.2 to happen with $g=1$, i.e. for the rank $1$ Drinfeld module $T\mapsto\alpha+\tau$, with $\alpha$ root of $\mathfrak{q}$ in ${\mathbb{F}}_{q^{m}}$.
Now derive the wanted parameters from Theorem 3.1. ∎